# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

### v3

@@ VII Conclusion
This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field static localization scenario. An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed. Extensive simulations have demonstrated that by optimizing the RIS phase shifts, the theoretical localization accuracy has been improved by two orders of magnitude. Moreover, the achievable localization accuracy can be lower than $\rm 10^{-2}\>m$. These findings corroborate the effectiveness of the proposed algorithms. To further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions with high computational complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This will be our future work.
@@ Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
$\frac{\partial\bm{\mu}[n]}{\partial d_{Ru}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial d_{Ru}}[\mathbf{S}]_{:,n},$ (55d)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{az}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{az}}[\mathbf{S}]_{:,n},$ (55e)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{el}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{el}}[\mathbf{S}]_{:,n},$ (55f)
@@
$\frac{\partial\bm{\mu}[n]^{*}}{\partial d_{Ru}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial d_{Ru}}\right)^{\rm H}\bm{\bar{w}},$ (56d)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{az}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{az}}\right)^{\rm H}\bm{\bar{w}},$ (56e)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{el}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{el}}\right)^{\rm H}\bm{\bar{w}}.$ (56f)

### v4

@@ VII Conclusion
This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field localization scenario. An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed. Extensive simulations have demonstrated that by optimizing the RIS phase shifts, the theoretical localization accuracy has been improved by two orders of magnitude. Moreover, the achievable localization accuracy can be lower than $\rm 10^{-2}\>m$. These findings corroborate the effectiveness of the proposed algorithms. To further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions with high computational complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This will be our future work.
@@ Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}[\mathbf{S}]_{:,n},$ (59d)
$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}[\mathbf{S}]_{:,n},$ (59e)
$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}[\mathbf{S}]_{:,n},$ (59f)
@@
$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\right)^{\rm H}\bm{\bar{w}},$ (60d)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}\right)^{\rm H}\bm{\bar{w}},$ (60e)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\right)^{\rm H}\bm{\bar{w}}.$ (60f)